Depreciation on a car can be determined by the formula V=C(1-r)^t , where V is the value of the car after t years, C is the original cost, and r the rate of depreciation. If a car's cost, when new, is $15,000, the rate of depreciation is 30%, and the value of the car now is $3,000, how old is the car to the nearest tenth of a year?
Answer by lwsshak3(11628)

Depreciation on a car can be determined by the formula V=C(1-r)^t , where V is the value of the car after t years, C is the original cost, and r the rate of depreciation. If a car's cost, when new, is $15,000, the rate of depreciation is 30%, and the value of the car now is $3,000, how old is the car to the nearest tenth of a year?
..
V=C(1-r)^t
3000=15000(1-.3)^t
3000=15000(.7)^t
3000/15000=(.7)^t
1/5 =(.7)^t
.2=.7^t
using logarithms to solve
log.2 = t*log.7
t=log.2/log.7
t=4.51 years
ans:
The car is 4.51 years old